OpenStudy (anonymous):

How do I Convert 15yd2 to square feet

OpenStudy (anonymous):

1 yd is 3ft so 1yd^2 is 9ft^2 so 15 yd^2 = 9 * 15 ft^2 = 135 ft^2

OpenStudy (anonymous):

A yard is equal to 3 feet. A yard^2 is therefore equal to 9 feet^2. 15*9=135 ft^2

OpenStudy (anonymous):

Thank You!